What is the vertex of the parabola: y= -2x^2+2x+4? When i solved it, i got the vertex as (-2,0) but all the online calculators say the vertex is (.5,4.5). Why is this? can you explain how to find the vertex?
Answer by stanbon(75887)

What is the vertex of the parabola: y= -2x^2+2x+4? When i solved it, i got the vertex as (-2,0) but all the online calculators say the vertex is (.5,4.5). Why is this? can you explain how to find the vertex?
-------------
The vertex of a parabola is always at x = -b/2a
In your problem a = -2 ; b = 2 ; c = 4
x = -b/2a = -2/2(-2) = 1/2 = 0.5
--------------------------------------------
Cheers,
Stan H.